find the sum of infinit series 12+22x+32x2+42x3+.................

Grade:12

1 Answers

PRAJVAL98
9 years ago
let its sum be S.
S=1+22x+32x2….n2xn-1=1+4x+9x2….…...n2xn-1
xS=x+4x2+9x3….(n-1)2xn-1+n2xn
S(1-x)=1+3x+5x2+7x3….-n2xn
 
use AGP formula to simplify it    LET THE TERM U GET AFTER SIMPLIFYING IT BE= $
 
S=($-n2xn)/(1-x)